# TPV: Parameter Perturbations Through the Lens of Test Prediction Variance

### v2

@@ 5.2 Empirical Evaluation of TPV Under Label Noise
Clean inputs are drawn i.i.d. from $\mathcal{N}(0,I)$ and targets follow a linear teacher $y=x^{\top}w_{\text{true}}$. Noisy targets are obtained by adding Gaussian noise with standard deviation $\sigma=0.01$. We train MLPs with widths $\{128,256,512,800,1024,1600\}$. Full details appear in Appendix F.2.1. Figure 5 and Figure 14 (Appendix) summarize the results.
@@
As the geometric regime changes with width, we again find that models with larger width have lower test loss and exhibit lower empirical TPV, and that training and test TPV estimates are similar in magnitude and trend (TPV stability). CIFAR-10 results and experiment details appear in Appendix F.2.2. A similar experiment on Imagenet where we track empirical TPV along with the validation accuracy is shown in Fig. 6 (see Appendix F.2.3 for details).
@@
5.3 Pruning
We evaluate whether the TPV-motivated pruning criterion (JBR) improves accuracy–compression tradeoffs relative to standard groupwise criteria. Following the OBC pruning protocol, we perform global channel pruning with iterative removal and recomputation of importance scores. We compare JBR against Jacobian, L1-norm, BatchNorm-scale, FPGM, WHC, Taylor, and Random. We prune two ImageNet models (ResNet-50, MobileNet-V2) at $50\%$ global sparsity and two CIFAR models (ResNet-56 on CIFAR-10, VGG-19-BN on CIFAR-100) at $90\%$ sparsity. Each model is pruned iteratively for 18 steps, averaging results over 5 independent runs. Results are shown in Fig. 9 and Fig. 10 in terms of MACs. Across all architectures, the JBR criterion consistently matches or exceeds the performance of all baselines. See Appendix F.3 for details.
6 Discussion and conclusion
We introduced TPV as a simple but powerful variance-based framework for probing the robustness of a trained neural network under specific noise sources. We proved that in overparameterized regime, training-set TPV tends to test-set TPV, irrespective of the model’s generalization performance, and empirically showed that this TPV stability holds remarkably well even in extremely narrow networks, as long as a sufficient number of samples is used, and the induced parameter perturbation covariance is small. This leads us to believe that Theorem 2.1 may be a sufficient condition under which TPV stability holds, but it may not be a necessary one.
@@ 6 Discussion and conclusion
A key advantage of using TPV over global variance in the global bias-variance decomposition is that TPV studies the properties of a specific trained model rather than all the possible models obtainable by a learning algorithm, which is also what we ultimately care about in real world– robustness of a trained model rather than the properties of the algorithm that trained it. Similar to how modern global bias-variance studies show that bias and variance typically tend to walk hand-in-hand in modern deep networks, we find that TPV estimates tend to correlate with standard test set loss across real world datasets and architectures.
@@
TPV can be used to measure the robustness of a trained model w.r.t. specific perturbation sources and therefore acts as a surgical tool to study generalization at a fine-grained level. In this paper, we showed that studying TPV for both SGD noise (Theorem 3.3) and quantization noise (Remark 3.4) leads to the wide minima hypothesis. This hypothesis is widely used as a tool for model selection and even guiding optimization (Foret et al., 2020). Under the TPV framework, we see that this hypothesis probes the robustness of a model under a specific class of parameter perturbations. Modeling label noise as a source of perturbation under the TPV framework leads to a different TPV estimate (Theorem 3.2) that depends on the spectrum of the output-parameter Jacobian and its alignment between training and test data distributions, and revealed a different practical approach to measuring TPV. We empirically found this object to also correlate with the standard test set loss for various datasets and models. One can similarly probe TPV under perturbations induced by a different noise source.
We believe TPV framework can serve as a foundation for future work on: (i) a deep empirical analysis that compares and contrasts a single model’s robustness w.r.t. different noise sources using the TPV framework; (ii) formalizing new noise sources other than the ones studied in this work (label noise, SGD noise, etc.); for instance, input perturbations (adversarial or domain shift); (iii) improving the algorithm for empirically estimating label noise TPV (see Appendix C.3); (iv) extending the TPV stability analysis to finite-width networks and TPV derivations to loss functions other than MSE.
@@ A.1 Assumptions
Since $H_{\mathrm{eff}}(w;X)$ and $G_{X}(w)$ share their nonzero spectrum, Assumption A.2 implies
$\|\,H_{\mathrm{eff}}(w^{\star};X)-H_{\mathrm{eff}}(w_{0};X)\,\|_{\mathrm{op}}~\leq~\varepsilon_{\mathrm{NTK}}.$ (22)
We assume $\|C\|_{F}<\infty$, which holds for any fixed finite covariance.
@@ A.3 Main Result
Theorem A.4 (TPV trace stability on finite train/test split).
Under Assumptions A.1 and A.2, and $C\succeq 0$, for the trained parameter $w^{\star}$,
$\bigl|\,\mathrm{TPV}(w^{\star};X_{\mathrm{tr}})-\mathrm{TPV}(w^{\star};X_{\mathrm{te}})\,\bigr|\leq(2\varepsilon_{\mathrm{NTK}}+o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1))\,\mathrm{Tr}(C).$
@@
Using (20) and symmetry of $H_{\mathrm{eff}}$ and $C$,
$\bigl|\mathrm{TPV}(w^{\star};X_{\mathrm{tr}})-\mathrm{TPV}(w^{\star};X_{\mathrm{te}})\bigr|=\bigl|\,\mathrm{Tr}\bigl(\bigl[H_{\mathrm{eff}}(w^{\star};X_{\mathrm{tr}})-H_{\mathrm{eff}}(w^{\star};X_{\mathrm{te}})\bigr]\,C\bigr)\bigr|.$
Now applying the following inequality for a PSD matrix $B$ and an arbitrary symmetric matrix AA,
$|\mathrm{Tr}(AB)|\leq\|A\|_{\mathrm{op}}\mathrm{Tr}(B),$
gives
$\bigl|\mathrm{TPV}(w^{\star};X_{\mathrm{tr}})-\mathrm{TPV}(w^{\star};X_{\mathrm{te}})\bigr|\leq\mathrm{Tr}(C)\,\|H_{\mathrm{eff}}(w^{\star};X_{\mathrm{tr}})-H_{\mathrm{eff}}(w^{\star};X_{\mathrm{te}})\|_{\mathrm{op}}.$
Using the triangle inequality and the operator-norm stability (22),
$\|H_{\mathrm{eff}}(w^{\star};X_{\mathrm{tr}})-H_{\mathrm{eff}}(w^{\star};X_{\mathrm{te}})\|_{\mathrm{op}}$
$\qquad\leq\|H_{\mathrm{eff}}(w^{\star};X_{\mathrm{tr}})-H_{\mathrm{eff}}(w_{0};X_{\mathrm{tr}})\|_{\mathrm{op}}+\|H_{\mathrm{eff}}(w_{0};X_{\mathrm{tr}})-H_{\mathrm{eff}}(w_{0};X_{\mathrm{te}})\|_{\mathrm{op}}$
$\qquad\quad+\|H_{\mathrm{eff}}(w_{0};X_{\mathrm{te}})-H_{\mathrm{eff}}(w^{\star};X_{\mathrm{te}})\|_{\mathrm{op}}$
$\qquad\leq 2\varepsilon_{\mathrm{NTK}}+o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1),$
where the $o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1)$ term is from (23). Thus,
$\bigl|\mathrm{TPV}(w^{\star};X_{\mathrm{tr}})-\mathrm{TPV}(w^{\star};X_{\mathrm{te}})\bigr|\leq(2\varepsilon_{\mathrm{NTK}}+o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1))\,\mathrm{Tr}(C).$
@@
Appendix B TPV for Linear Regression under Label Noise: Benign Overfitting
@@
Appendix C TPV for Non-Linear Regression Under Label Noise
C.1 Proof
@@
C.2 Useful Regime and Pathological Case Analysis
@@
C.3 Practical Computation of the Label–Noise TPV Quantity
C.3.1 Idealized Linearized Response Under Label Noise
@@
C.3.2 Dual Formulation and Ideal Numerical Computation
@@
C.3.3 Empirical Observation: Deep–Network Jacobians Are Extremely Ill–Conditioned
@@
C.3.4 Practical Resolution: Local SGD With Proximity Regularization
@@
Appendix D TPV for SGD Stationary Noise
@@
D.1 Relation Between Gradient Covariance Matrix and Hessian near Minima
@@
D.2 Relation Between $H_{\mathrm{eff}}$ And Hessian Near Minima
@@
D.3 SGD Late Dynamics and Relation Between TPV and Hessian
@@
where $\nabla_{w}^{2}L(w)$ is the Hessian. We now use the result from appendix D.2 stating $H_{\mathrm{eff}}\approx\nabla_{w}^{2}L(w)$. Thus,
@@
Finally, we use the result from Appendix D.1 and D.2, and assuming that the Hessian is stable around the minimum $w^{\star}$, which lies at the center of the SGD stationary dynamics, we have that TPV under SGD noise is given by,
@@
Appendix E TPV for Parameter Quantization Noise
TPV is given by $\mathrm{Tr}(H_{\mathrm{eff}}C)$. We show in Appendix D.2 that $H_{\mathrm{eff}}(w)\approx\nabla_{w}^{2}L(w)$ (Hessian) near minima $w^{\star}$. Under the quantization model above, the parameter perturbation covariance is $C_{\mathrm{quant}}\approx\frac{\delta^{2}}{12}I_{p}$, where $I_{p}$ is the identity matrix (thus obeying the TPV trace stability requirement), which proves the claim.
Appendix F Experiments
@@
F.1 Details of Experiments in Section 5.1
F.1.1 Synthetic Data Experiment:
@@ 1. Vary Width Experiment Details:
Label-noise mechanism (logit noise). For the label-noise experiments, we retrain from the fixed initialization $w^{\star}$ for $R=5$ independent runs. Each run injects additive i.i.d. Gaussian noise $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ with $\sigma\in\{0.05,0.1\}$ to the clean labels. Each model is then trained for $50$ epochs using mini-batch SGD MSE regression on the noisy logits with momentum $0.9$, learning rate $10^{-4}$, batch size $256$, and 0 weight decay. Mini-batch shuffling is turned off and a randomness seed is used so that each run sees the same sequence of mini-batches in order to avoid randomness due to SGD and focus only on randomness due to label noise. Also, we train the models in Pytorch eval mode so that modules like batch norm do not make output logits batch dependent. See Appendix C.3.4 for more details. We record the train/test CE (for generalization gap) and empirical train/test TPV using all these runs.

### v3

@@ 5.2 Empirical Evaluation of TPV Under Label Noise
Clean inputs are drawn i.i.d. from $\mathcal{N}(0,I)$ and targets follow a linear teacher $y=x^{\top}w_{\text{true}}$. Noisy targets are obtained by adding Gaussian noise with $\sigma=0.01$. We train MLPs with widths $\{128,256,512,800,1024,1600\}$. Full details appear in Appendix G.2.1. Fig. 4 and Fig. 12 summarize the results.
@@
As the geometric regime changes with width, we again find that models with larger width have lower test loss and exhibit lower empirical TPV, and that training and test TPV estimates are similar in magnitude and trend (TPV stability). CIFAR-10 results and experiment details appear in Appendix G.2.2. A similar experiment on Imagenet where we track empirical TPV along with the validation accuracy is shown in Fig. 5 (see Appendix G.2.3 for details).
@@
6 Conclusion and future work
We introduced TPV as a simple yet powerful variance-based framework for probing the robustness of a trained neural network under specific noise sources. The TPV stability theorem formalizes a sufficient mechanism—i.i.d. sampling and isotropic perturbations under NTK stability in the overparameterized regime—under which TPV stability holds, while our experiments suggest that TPV stability holds more broadly, including in narrow networks and for structured perturbations beyond this setting. This gap highlights TPV stability as an intrinsic empirical property of trained networks and motivates developing theory beyond NTK-based and isotropic assumptions.
@@ 6 Conclusion and future work
A key advantage of TPV over global variance in the classical bias–variance decomposition is that TPV characterizes the behavior of a specific trained model rather than all the models obtainable by a learning algorithm, which is more directly aligned with real-world robustness concerns. We further find that TPV estimates tend to correlate with standard test-set loss across datasets and architectures, mirroring empirical observations in modern bias–variance studies.
@@
TPV can be used to measure the robustness of a trained model with respect to specific perturbation sources, providing a fine-grained tool for studying generalization. In this paper, we showed that TPV analyses under SGD noise (Theorem 3.3) and quantization noise (Remark 3.4) recover the wide-minima hypothesis, which is widely used as a tool for model selection and even guiding optimization (Foret et al., 2020), while modeling label noise as a perturbation leads to a distinct TPV estimate (Theorem 3.2) governed by the spectrum and alignment of the output-parameter Jacobian between training and test distributions. We empirically found this quantity to also correlate with test-set loss across datasets and models.
We believe the TPV framework provides a foundation for future work on: (i) comparative empirical studies of robustness under different noise sources; (ii) formalizing additional perturbation models (e.g. input perturbations and distribution shift); (iii) improving empirical estimation of label-noise TPV (Appendix D.3); and (iv) relaxing the assumptions in the TPV stability analysis and extending it beyond the overparameterized regime and beyond MSE loss.
@@ A.1 Assumptions
Assumption A.3 (Isotropic Covariance).
The perturbation covariance matrix is isotropic, i.e., $C=\sigma^{2}\mathbf{I}_{p}$, where $p$ is the number of parameters and $\sigma$ is a scalar.
@@ A.3 Main Result
We now state and prove a finite-sample TPV trace stability result.
Theorem A.5 (TPV trace stability under isotropic perturbations).
Under Assumptions A.1, A.2, and A.3, we have
$\bigl|\mathrm{TPV}(w^{\star};X_{\mathrm{tr}})-\mathrm{TPV}(w^{\star};X_{\mathrm{te}})\bigr|~\leq~\mathrm{Tr}(C)\left(\,\frac{(n_{\mathrm{tr}}+n_{\mathrm{te}})}{p}\,\varepsilon_{\mathrm{NTK}}~+~o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1)\right),$ (19)
@@
By Assumption A.3, $C=\sigma^{2}\mathbf{I}_{p}$, so
$\mathrm{TPV}(w;X)$ $=\mathrm{Tr}\!\left(H_{\mathrm{eff}}(w;X)\,\sigma^{2}\mathbf{I}_{p}\right)~=~\sigma^{2}\,\mathrm{Tr}\!\left(H_{\mathrm{eff}}(w;X)\right)$
$=\sigma^{2}\,\mathrm{Tr}\!\left(\frac{1}{n_{X}}J_{X}(w)^{\top}J_{X}(w)\right)~=~\sigma^{2}\,\mathrm{Tr}\!\left(\frac{1}{n_{X}}J_{X}(w)J_{X}(w)^{\top}\right)~=~\sigma^{2}\,\mathrm{Tr}\!\left(G_{X}(w)\right).$ (20)
We begin with a triangle inequality decomposition:
$\bigl|\mathrm{TPV}(w^{\star};X_{\mathrm{tr}})-\mathrm{TPV}(w^{\star};X_{\mathrm{te}})\bigr|$
$\qquad\leq\bigl|\mathrm{TPV}(w^{\star};X_{\mathrm{tr}})-\mathrm{TPV}(w_{0};X_{\mathrm{tr}})\bigr|+\bigl|\mathrm{TPV}(w_{0};X_{\mathrm{tr}})-\mathrm{TPV}(w_{0};X_{\mathrm{te}})\bigr|+\bigl|\mathrm{TPV}(w_{0};X_{\mathrm{te}})-\mathrm{TPV}(w^{\star};X_{\mathrm{te}})\bigr|.$ (21)
Step 1: Controlling the train/test drift from initialization via NTK stability:
Fix a dataset $X$ of size $n_{X}$. Using (20),
$\bigl|\mathrm{TPV}(w^{\star};X)-\mathrm{TPV}(w_{0};X)\bigr|=\sigma^{2}\,\bigl|\mathrm{Tr}\bigl(G_{X}(w^{\star})-G_{X}(w_{0})\bigr)\bigr|.$
For any matrix $A\in\mathbb{R}^{n_{X}\times n_{X}}$, $|\mathrm{Tr}(A)|\leq\mathrm{rank}(A)\,\|A\|_{\mathrm{op}}\leq n_{X}\|A\|_{\mathrm{op}}$, hence
$\bigl|\mathrm{TPV}(w^{\star};X)-\mathrm{TPV}(w_{0};X)\bigr|\leq\sigma^{2}\,n_{X}\,\|G_{X}(w^{\star})-G_{X}(w_{0})\|_{\mathrm{op}}.$
Applying Assumption A.2 yields, for $X\in\{X_{\mathrm{tr}},X_{\mathrm{te}}\}$,
$\bigl|\mathrm{TPV}(w^{\star};X)-\mathrm{TPV}(w_{0};X)\bigr|\leq\sigma^{2}\,n_{X}\,\varepsilon_{\mathrm{NTK}}.$ (22)
Step 2: Controlling the initialization train–test gap via LLN:
Using (16) and Assumption A.3,
$\mathrm{TPV}(w_{0};X)=\sigma^{2}\,\mathrm{Tr}\!\left(H_{\mathrm{eff}}(w_{0};X)\right).$
Therefore,
$\bigl|\mathrm{TPV}(w_{0};X_{\mathrm{tr}})-\mathrm{TPV}(w_{0};X_{\mathrm{te}})\bigr|=\sigma^{2}\,\bigl|\mathrm{Tr}\!\left(H_{\mathrm{eff}}(w_{0};X_{\mathrm{tr}})-H_{\mathrm{eff}}(w_{0};X_{\mathrm{te}})\right)\bigr|.$
For symmetric $B\in\mathbb{R}^{p\times p}$, $|\mathrm{Tr}(B)|\leq p\|B\|_{\mathrm{op}}$, so
$\bigl|\mathrm{TPV}(w_{0};X_{\mathrm{tr}})-\mathrm{TPV}(w_{0};X_{\mathrm{te}})\bigr|\leq\sigma^{2}\,p\,\|H_{\mathrm{eff}}(w_{0};X_{\mathrm{tr}})-H_{\mathrm{eff}}(w_{0};X_{\mathrm{te}})\|_{\mathrm{op}}.$
Applying Lemma A.4 gives
$\bigl|\mathrm{TPV}(w_{0};X_{\mathrm{tr}})-\mathrm{TPV}(w_{0};X_{\mathrm{te}})\bigr|\leq\sigma^{2}\,p\cdot o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1).$ (23)
Step 3: Combine:
Plugging (22) (for $X_{\mathrm{tr}}$ and $X_{\mathrm{te}}$) and (23) into (21) yields
$\bigl|\mathrm{TPV}(w^{\star};X_{\mathrm{tr}})-\mathrm{TPV}(w^{\star};X_{\mathrm{te}})\bigr|\leq\sigma^{2}\,n_{\mathrm{tr}}\,\varepsilon_{\mathrm{NTK}}+\sigma^{2}\,p\cdot o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1)+\sigma^{2}\,n_{\mathrm{te}}\,\varepsilon_{\mathrm{NTK}},$
which is exactly (19). ∎
Appendix B TPV for Scalar and Vector Output Models
We clarify the definition of the $H_{\mathrm{eff}}$ and Test Prediction Variance (TPV) for both scalar- and vector-output models.
Scalar-output case. Let $f_{w}(x)\in\mathbb{R}$ be a scalar-output model with parameters $w\in\mathbb{R}^{p}$, and define the Jacobian $J(x;w)=\nabla_{w}f_{w}(x)\in\mathbb{R}^{1\times p}$. For a dataset $X=\{x_{i}\}_{i=1}^{n}$, $H_{\mathrm{eff}}$ is
$H_{\mathrm{eff}}(w;X)\;=\;\frac{1}{n}\sum_{i=1}^{n}J(x_{i};w)^{\top}J(x_{i};w)\;=\;\frac{1}{n}J_{X}(w)^{\top}J_{X}(w),$
where $J_{X}(w)\in\mathbb{R}^{n\times p}$ stacks the per-sample Jacobians. For a zero-mean parameter perturbation $\delta w$ with covariance $C$, the (first-order) TPV is
$\mathrm{TPV}(X)\;=\;\frac{1}{n}\mathbb{E}_{\delta w}\!\left[\|f_{w+\delta w}(X)-f_{w}(X)\|_{2}^{2}\right]\;\approx\;\mathrm{Tr}\!\left(H_{\mathrm{eff}}(w;X)\,C\right).$
Vector-output case. Let $f_{w}(x)\in\mathbb{R}^{K}$ (e.g., logits) and $J(x;w)=\nabla_{w}f_{w}(x)\in\mathbb{R}^{K\times p}$. $H_{\mathrm{eff}}$ is defined as
$H_{\mathrm{eff}}(w;X)\;=\;\frac{1}{n}\sum_{i=1}^{n}J(x_{i};w)^{\top}J(x_{i};w),$
which is again a $p\times p$ matrix. The definition of TPV now becomes,
$\mathrm{TPV}(X)\;=\;\frac{1}{n}\mathbb{E}_{\delta w}\!\left[\sum_{i=1}^{n}\|f_{w+\delta w}(x_{i})-f_{w}(x_{i})\|_{2}^{2}\right]\;\approx\;\mathrm{Tr}\!\left(H_{\mathrm{eff}}(w;X)\,C\right).$
In particular, for isotropic perturbations $C=\sigma^{2}I$, $\mathrm{TPV}(X)=\sigma^{2}\,\mathrm{Tr}(H_{\mathrm{eff}}(w;X))$ in both cases.
@@ Appendix B TPV for Scalar and Vector Output Models
Appendix C TPV for Linear Regression under Label Noise: Benign Overfitting
@@
Appendix D TPV for Non-Linear Regression Under Label Noise
D.1 Proof
@@
D.2 Useful Regime and Pathological Case Analysis
@@
D.3 Practical Computation of the Label–Noise TPV Quantity
D.3.1 Idealized Linearized Response Under Label Noise
@@
D.3.2 Dual Formulation and Ideal Numerical Computation
@@
D.3.3 Empirical Observation: Deep–Network Jacobians Are Extremely Ill–Conditioned
@@
D.3.4 Practical Resolution: Local SGD With Proximity Regularization
@@
Appendix E TPV for SGD Stationary Noise
@@
E.1 Relation Between Gradient Covariance Matrix and Hessian near Minima
@@
E.2 Relation Between $H_{\mathrm{eff}}$ And Hessian Near Minima
@@
E.3 SGD Late Dynamics and Relation Between TPV and Hessian
@@
where $\nabla_{w}^{2}L(w)$ is the Hessian. We now use the result from appendix E.2 stating $H_{\mathrm{eff}}\approx\nabla_{w}^{2}L(w)$. Thus,
@@
Finally, we use the result from Appendix E.1 and E.2, and assuming that the Hessian is stable around the minimum $w^{\star}$, which lies at the center of the SGD stationary dynamics, we have that TPV under SGD noise is given by,
@@
Appendix F TPV for Parameter Quantization Noise
TPV is given by $\mathrm{Tr}(H_{\mathrm{eff}}C)$. We show in Appendix E.2 that $H_{\mathrm{eff}}(w)\approx\nabla_{w}^{2}L(w)$ (Hessian) near minimum $w^{\star}$. Under the quantization model above, the parameter perturbation covariance is $C_{\mathrm{quant}}\approx\frac{\delta^{2}}{12}I_{p}$, where $I_{p}$ is the identity matrix (thus obeying the TPV trace stability requirement), which proves the claim.
Appendix G Experiments
@@
G.1 Details of Experiments in Section 5.1
G.1.1 Synthetic Data Experiment:
@@ 1. Vary Width Experiment Details:
Label-noise mechanism (logit noise). For the label-noise experiments, we retrain from the fixed initialization $w^{\star}$ for $R=5$ independent runs. Each run injects additive i.i.d. Gaussian noise $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ with $\sigma\in\{0.05,0.1\}$ to the clean labels. Each model is then trained for $50$ epochs using mini-batch SGD MSE regression on the noisy logits with momentum $0.9$, learning rate $10^{-4}$, batch size $256$, and 0 weight decay. Mini-batch shuffling is turned off and a randomness seed is used so that each run sees the same sequence of mini-batches in order to avoid randomness due to SGD and focus only on randomness due to label noise. Also, we train the models in Pytorch eval mode so that modules like batch norm do not make output logits batch dependent. See Appendix D.3.4 for more details. We record the train/test CE (for generalization gap) and empirical train/test TPV using all these runs.